### v2

# Geographic Variation in

Abstract
Using complete-count register data spanning three generations, we document spatial patterns in inter- and multigenerational mobility in Sweden. Across municipalities, grandfather-child correlations in education or earnings tend to be larger than the square of the parent-child correlations, suggesting that the latter understate status transmission in the long run. Yet, conventional parent-child correlations capture regional differences in long-run transmission and therefore remain useful for comparative purposes. We further find that the within-country association between mobility and income inequality (the “Great Gatsby Curve”) is at least as strong in the multi- as in the intergenerational case. Interpreting those patterns through the lens of a latent factor model, we find that regional differences in mobility primarily reflect variation in the transmission of latent advantages, rather than in how those advantages translate into observed outcomes.
JEL Codes: J62, R00
Keywords: The geography of intergenerational mobility, multigenerational mobility, income inequality
1 Introduction
An increasingly popular approach to study intergenerational mobility is to compare mobility rates across regions. Theoretical models with social or spatial dimensions (e.g. Durlauf1996 and Benabou1996a) suggest that income persists across generations partly because parents shape their children’s social environment, such as schools or neighborhoods. Empirically, regional variations can be used to identify causal effects; for example, Holmlund2006 and PekkarinenUusitaloKerr2009 exploit the staggered implementation of an educational reform to estimate its impact on intergenerational mobility. Following CHKS_Where_QJE, some studies invert this research design: rather than focusing on a specific policy or event, they first map regional variation in intergenerational mobility and then search for regional correlates that could explain this variation. [^1] Examples include ConnollyCorakHaeck2019JOLE, ConnollyHaeckLapierre2019 and Corak2017Landscapes for Canada, DeutscherMazumder2020Labour for Australia, Heidrich:2017aa and Branden2019 for Sweden, Risa:2019aa and butikofer2022breaking for Norway, ERIKSEN2020109024 for Denmark, acciari2022and for Italy, or BellBlundellMachin2023 for the UK.
In this study, we extend this “regional approach” by considering not only inter- but also multigenerational associations. Such associations are interesting both from a descriptive and a causal perspective. Recent studies indicate that socioeconomic disparities are more persistent than extrapolations from the available parent-child correlations would suggest (AndersonSheppardMonden2017; Solon_2018; stuhler2024multigenerational). This contrast between inter- and multigenerational correlations is in turn informative about the statistical properties and potential mechanisms of mobility processes (e.g., ColladoOrtunoStuhlerKinship). By adopting such multigenerational perspective, we aim to better understand the causes and implications of regional differences in mobility.
We make four related contributions: First, we describe spatial patterns in multigenerational mobility in education and earnings across Swedish municipalities. Second, we study whether those patterns are robust to the choice of mobility measure and the quality of the underlying microdata. Third, we study whether earnings inequality and mobility are related, and whether the strength of this association varies across different mobility statistics. Finally, we compare inter- and multigenerational correlations through the lens of a latent transmission model to gain a deeper insight into how mobility varies across regions.
We first report regional estimates of educational and earnings mobility across three generations for cohorts born in the 1980s across 290 municipalities. Municipalities are key geographical entities in the Swedish context, as many public goods and schooling are provided at this level, and often correspond to the relevant labor market of their inhabitants. Our evidence here is complementary to evidence provided by Heidrich:2017aa and Branden2019, who estimate income mobility for earlier birth cohorts and a broader definition of local labor markets or commuting zones.
We then study whether such regional rankings are sensitive to the choice of outcome (education vs. earnings), statistic (e.g., regression vs. correlation coefficients), time frame (inter- vs. multigenerational), or lineage (paternal vs. maternal line). One particular question is whether summary statistics based on education – which can be estimated in most settings – can serve as a substitute for more data-demanding measures based on income. This part of our analysis is complementary to recent work by deutscher2023measuring, who compare the ranking of Australian regions across different measures of income mobility.
We find that regional rankings based on inter- and multigenerational measures are similar. Municipalities with low intergenerational mobility in education (high parent-child correlations) also tend to have low multigenerational correlations (high grandparent-child correlations). In contrast, the association between education- and earnings-based mobility measures is weak. While this observation contrasts with cross-national comparisons, in which educational and income mobility are found to be closely related (Blanden2011Survey; Stuhler2018JRC), it is reminiscent of the sociological discourse on the “mobility paradox”, with some countries showing diverging patterns of income compared to educational or occupational mobility (e.g. breen2016much; karlson2021denmark).
We next study whether multigenerational mobility and cross-sectional inequality are systematically related across regions. A well-documented pattern in international comparisons shows a negative association between income mobility and inequality (Blanden2011Survey, CorakJEP2013). We find that the same relationship holds across municipalities within Sweden, not only for inter- but also for multigenerational mobility. In line with evidence by Branden2019, who analyzed father-child mobility for earlier Swedish cohorts, we observe that municipalities that are characterized by high income inequality in the parent’s generation also tend to exhibit low mobility. However, the strength of this relationship varies considerably with the choice of mobility statistic. Educational mobility is more strongly correlated with inequality than income mobility, possibly due to greater measurement error in income data. Interestingly, the relationship with inequality is particularly pronounced for multigenerational correlations in education.
Finally, we interpret regional variations in inter- and multigenerational correlations through the lens of a latent factor model that allows for the transmission of unobserved advantages from parents to children. Estimating its parameters separately for each municipality, we find that regional variation in intergenerational mobility is primarily due to variation in the transmission process (i.e., the “transferability” of latent advantages from one generation to the next), rather than regional variation in how latent advantages map into observed advantages in education or earnings (i.e., the “returns” to those latent advantages). One implication of this finding is that regional differences in parent-child correlations may understate the differences in long-run mobility across regions.
One important consideration in our analysis is the role of sampling error, which will be more pronounced in regional compared to national-level analyses. While we briefly explore its influence, a more systematic examination would be useful for future research. Multigenerational comparisons across regions may also be sensitive to shifts in marginal distributions across generations, in particular if those shifts occur at different times across regions. While these are important caveats, our estimates illustrate the potential benefits of integrating multigenerational and “regional” approaches to studying social mobility.
2 Data
Our main data sources are administrative registers that cover the universe of Swedish citizens aged 16-64 at any point between the years 1960-2020 and gross labor earnings from tax records (reported by employers) for the period 1968-2020. Using multigenerational registers, we link children born in Sweden 1932 or later, as well as immigrants who have been residents of Sweden at some point since 1961, to their biological parents. Other administrative registers and the bidecennial censuses 1960-1990 provide information on education, place of birth and current residence, and other individual characteristics. In particular, the education register contains data on highest educational attainment and field of education for the population alive in 1970 or later (and born in 1911 or later).
2.1 Samples and outcomes
We focus on cohorts born between 1981 and 1989, as these cohorts have had sufficient time to complete their education and enter the labor market, and we are able to merge educational or income information for their parents (intergenerational analysis) and in the majority of cases also their grandparents (multigenerational analysis). [^2] We are able to identify parents and grandparents if the parent was born in 1932 or later (or 1961 or later if the parent immigrated to Sweden). As our earliest child cohort is born in 1981, this restriction is not binding for the matching of parents (98.3 percent and 96.7 percent match rates for mothers and fathers, respectively) but it slightly reduces the match rate of grandparents (88 percent matched to at least one grandparent; 90 percent are matched to the paternal grandfather). When also requiring non-missing education or income the match rates and sample sizes decrease (see further below). We allocate families to regions based on the child’s municipality of residence when the child was aged 16. Those who lack information on place of residence at age 16 (primarily due to living abroad) are excluded from all analyses. In a robustness test, we instead allocate families according to the mother’s location when they were young. [^3] For each mother, we use the earliest municipality that we find in our data between 1976 and 1994. We do not use earlier records as the municipality codes were less comparable. We find a location for more than 96% of mothers in our sample, usually in 1976. As the average year of birth of sampled mothers is 1957, their locations are measured around age 20, typically during adolescence or in the early 20s.
Our primary outcome measures are years of schooling and labor earnings. To construct years of schooling, we consider seven different educational categories that have long been used by Statistics Sweden, and recode them to years of education following conventional procedures. [^4] Specifically, we distinguish short compulsory education in the old system (7 years of schooling); compulsory education (9 years); some secondary education (10.5 years); high-school degree (12 years); some college or post-secondary vocational education (14 years); college degree (16 years); PhD or equivalent (20 years). We lack education data for those (grandparents) born before 1911.
Our measure of labor earnings includes the annual sum (before taxes and transfers) of all earnings according to employer-reported tax and earnings statements to the national tax agency, which captures all types of labor compensation subject to payroll taxes. [^5] We observe labor earnings (including income from self-employment) for all residents in 1968, 1970, 1971, 1973, 1975, 1976, 1979, 1980, 1982 and annually for 1985-2020. We adjust annual incomes for inflation using the CPI. While it would have been interesting to use disposable income, the information on disposable income does not go as far back in time. The measure is not censored or top-coded, and includes bonus payments. For our purposes, it is crucial to use earnings measures that are as comparable as possible over generations, as our estimates would be distorted if the influence of lifecycle and attenuation biases varies across generations. Given the structure of our data set, with earnings observed over a given time period and different generations (naturally) born several decades apart, it is not possible to measure earnings over the same age ranges for all generations. We therefore use a Mincer-type approach to predict prime-age earnings for all generations, which accounts for differences in income growth across education groups and differences in income levels across cohorts. [^6] While this approach improves our earnings measures, they are still unlikely to be fully comparable across generations. Specifically, we do not account for the observation that age-education-earnings profiles may differ across cohorts (heckman2003fifty) or be steeper for children from parents with higher socioeconomic status (MelloNybomStuhlerLifecycle). However, we do allow for individual intercepts, thereby capturing general shifts in earnings levels across generations. Table A.1 in the Appendix compares area-level inter- and multigenerational earnings rank slopes for our prediction-based measure with more standard measures based on observed incomes. While the dispersion across municipalities is largely similar across measures, the prediction-based measure yields larger coefficients with magnitudes that are more in line with recent evidence for Sweden.
In a first step, we use a panel of all individuals in the labor force aged 25-63 over the years 1968-2020 with annual earnings corresponding to at least 25 percent of the male median earnings in a given year. [^7] We do not include individuals born before 1915 in the income analyses, as we want to avoid using unrepresentative income observations too close to retirement age. We then regress log annual earnings on individual fixed effects and gender-by-education groups separately interacted with quadratic polynomials in age and year. We use separate indicators for the seven levels of educational attainment (described above) but also allocate those without information on educational attainment to a separate “missing education” category. Lastly, we use the resulting estimates to predict for all individuals their log earnings at age 40 and then link this prediction to all members of the three generations of our multigenerational sample. [^8] We bottom code predicted annual earnings to SEK1000 (roughly 90 USD) at age 40 to reduce the influence of extreme negative outliers. For the child generation, because we regularly consider pooled analyses of sons and daughters, we also demean predicted log earnings by gender. However, this demeaning has generally very little influence on our results. Some parts of our analysis focus on earnings ranks; we rank the predicted earnings separately by relationship type (child, father, mother, paternal grandfather, etc.) and the child’s year of birth.
Table 1 shows our samples and outcome means. For both education and income, respectively, we work both with an “unbalanced” sample with non-missing outcomes for the child and the father, but possibly missing outcomes for the grandfather, and a “balanced” sample with non-missing outcomes for child, father and grandfather. We observe father’s education (earnings) for 96 (95) percent of those in the child cohorts with non-missing education (earnings). Among those with non-missing education (earnings) for both child and father, we also observe the same outcome for the paternal grandfather in about 70 (78) percent of cases. [^9] While education is observed slightly more often than (predicted) earnings for children and fathers, the opposite is true for grandparents, since we allow missing education (allocated to a separate category) in the earnings prediction. The balanced sample is therefore slightly larger for earnings than for education. For some analyses we consider other relatives, such as mothers, grandmothers, maternal grandfathers, or average across all relatives by generation.
Table 1 also reports mean birth years across generations (for those in the education samples). These statistics (for the unbalanced sample) imply that the mean age at birth of the child of fathers and mothers are 31 and 28, respectively. For grandfathers and grandmothers, their mean ages at birth of their grandchild are 62 and 59, respectively. These numbers are very similar to what we find in population-wide registers for the same child cohorts, before adding any further sampling restrictions.
There are 290 regions (municipalities) included in our analyses, with the smallest one containing 263 observations and the largest region containing 56,969 observations in our unbalanced education sample. Figure A.1 in the Appendix shows the distribution of individual observations across municipalities ordered from the smallest to largest, separately for our different analysis samples. The figure shows that about 20 percent of observations live in the 150 smallest observations, whereas about 50 percent of observations live in the 40 largest municipalities. Less than 5 percent of the observations in the sample come from municipalities with a sample size lower than 1,000. However, to reduce the influence of sampling error (see next section), we regularly weight regressions by the number of parent-child pairs per region. Some analyses are also restricted to larger municipalities with more precisely estimated statistics.
Table 1: Samples and Descriptive Statistics
Year of birth Years of education Log earnings
Unbal. Balanced Unbal. Balanced Unbal. Balanced
Observations 923,061 645,806 923,061 645,806 898,783 701,358
Means:
  Child 1985 1985 13.45 13.48 12.87 12.89
  Father 1954 1955 11.69 11.73 12.42 12.50
  Mother 1957 1958 12.17 12.29 12.07 12.12
  Grandfather (pat.) 1923 1925 9.17 9.17 12.47 12.48
  Grandmother (pat.) 1926 1928 8.79 8.85 11.41 11.41
Note: Sample size and means for each sample. The number of observations correspond to the number of distinct individuals in the child generation (born 1981-1989). The unbalanced samples have non-missing information on the child’s and father’s education (column 3) or log earnings (column 5). The balanced samples have non-missing information on the child’s, father’s, and grandfather’s education or log earnings. Before considering parental or grandparental information, there are 961,518 children in our cohorts with non-missing education and 941,157 with non-missing log income.
2.2 Mobility statistics
We consider a range of standard mobility statistics, such as the intergenerational (or multigenerational) regression coefficient,
$y_{ir}=\alpha_{r}+\beta_{r}y_{i}^{p}+\varepsilon_{ir},$ (1)
where $y_{ir}$ is one of the outcomes described above for individual $i$ in region (municipality) $r$ and $y_{i}^{p}$ is the corresponding outcome for a parent or grandparent. The slope $\beta_{r}$ measures “relative” mobility, comparing socioeconomic status relative to the status of others in the same generation. For education, we also consider the corresponding Pearson correlation coefficient, which abstracts from shifts in the variance between the two generations within the region. For our analysis of earnings mobility, we consider either log earnings (in which case the slope coefficient $\beta_{r}$ can be interpreted as the intergenerational elasticity, i.e. the IGE) or earnings ranks (in which case $\beta_{r}$ can be referred to as the rank slope). Ranks are defined on the national level, separately within each child cohort. [^10] Based on the estimated intercept $\hat{\alpha}_{r}$ and the slope $\hat{\beta}_{r}$ we further construct the expected “absolute upward mobility” of children born to parents at the 25th percentile (i.e., for $y_{i}^{p}=0.25$), as defined in CHKS_Where_QJE. Our primary focus is on education correlations or the earnings rank slope, and we focus only occasionally on the IGE and the measure of upward mobility.
3 Geographical Variation in Mobility
We first document that inter- and multigenerational mobility varies across municipalities within Sweden. To illustrate the extent of variation in educational mobility, Figure 1 plots the density of the correlation in years of schooling across 290 municipalities, separately for father-child, mother-child, grandfather-child and grandmother-child pairs (pooling sons and daughters). The within-municipality parent-child correlations are clustered around a (weighted) mean of 0.30, but vary from 0.09 to 0.42 (see also Table 2). The three-generation correlations are clustered around a mean of 0.11, again with large variation across municipalities. The means of both the two- and three-generation correlations are about 5-10% lower in the unweighted case.
Figure 1: Inter-and multigenerational correlations in education in Sweden
Notes: Density of the estimated intergenerational correlations in years of schooling for parent-child and grandparent-child pairs across 290 Swedish municipalities (bandwidth: 0.01), based on birth cohorts 1981-99. Observations are weighted by the number of parent-child pairs.
As daughters’ education resembles their mother’s more than their father’s education (the relations are more symmetric for sons), mother-child correlations tend to be marginally larger than the corresponding father-child correlations (0.31 vs 0.30); see Appendix Table A.2 for separate statistics by child gender. The inverse pattern applies for grandparent correlations, which are systematically larger for grandfather-child than for grandmother-child pairs (0.12 vs 0.09). [^11] These opposing gender patterns illustrate that the intergenerational transmission process may vary across generations, which constitutes a challenge for attempts to estimate transmission models from multigenerational data (NybomStuhler2019). Figure A.2 in the Appendix shows the corresponding distributions for earnings rank slope coefficients. In that case, the father-child coefficients are larger than those for mother-child pairs, reflecting that mothers’ earnings tend to be poorer proxies of underlying social status. Nevertheless, the differences between the two are quite small, reflecting that the mothers in our sample were already participating on the labor market to a comparatively large degree. Moreover, the grandfather-child coefficients follow a strikingly similar distribution for paternal and maternal-side grandfathers.
3.1 Comparing inter- and multigenerational mobility
Parent-child correlations as plotted in Figure 1 may understate the extent to which socio-economic advantages are transmitted from one generation to the next, because educational attainment is just one aspect of a person’s overall socio-economic status. Indeed, studies that combine multiple status proxies (VostersNybom2017; BlundellRisa2018aa), that integrate information from grandparents or the extended family (ECOJ:ECOJ12453; colagrossi2020like; Adermonetal2016; ColladoOrtunoStuhlerKinship), or that consider long-run mobility across many generations (Clark2014book; BaroneMocetti2016) tend to find higher persistence.
We illustrate this argument in two ways. First, we use generational averages rather than individual outcomes of a single parent or grandparent. In particular, we compute mean outcomes for both parents and all four grandparents, forming generation-specific means of either schooling or earnings ranks. By averaging over both parents, or all grandparents, these mean measures might provide a better proxy for socioeconomic status of a family. They might also reduce the influence of measurement and sampling error. [^12] Previous studies that construct averages across several family members of the same generation or even across the entire family tree include thaning2020end, Adermonetal2016, and hallsten2023shadow.
Table 2 shows summary statistics of inter- and multigenerational correlations for different lineages and measures across municipalities, either unweighted (columns 1 and 2) or weighted by municipality-specific sample size (columns 3 and 4). For the weighted case, the parent-child correlation in schooling increases slightly once the schooling of both parents are taken into account (from 0.30-0.31 to 0.36). A similar but stronger pattern is found for earnings ranks, with the mean coefficients increasing from about 0.20-0.22 to above 0.30.131313Note that since for earnings we report regression rather than correlation coefficients, this increase in magnitudes partly stems from a change in the relative variances of the dependent and independent variables. The grandparent-child correlations show similar tendencies: multigenerational correlations in schooling increase from around 0.09 (grandmothers) or 0.12 (grandfathers) to almost 0.15 when we average the schooling of all four grandparents. For earnings ranks, the correlations increase from around 0.09-0.10 for individual grandfathers to above 0.14 when averaging over all grandparents. The patterns are very similar for the unweighted case, though the estimates tend to be slightly lower.
As already alluded to, the correlations in Table 2 also vary with the gender of the child (see Table A.2 in the Appendix); for example, the (weighted) father-son correlation in years of schooling is on average about 0.33, while the father-daughter correlation is slightly lower at 0.27; and this difference is statistically significant ($p<0.01$). While the pattern flips when considering mothers only, the correlations for sons are slightly larger once the schooling of both parents are taken into account. Overall, the relative differences along gender-specific lines should however be regarded as fairly moderate. Sons also appear slightly less multigenerationally mobile; the schooling correlations tend to be larger for sons irrespectively of whether we consider individual grandparents or jointly take all grandparents into account (0.16 for sons vs. 0.13 for daughters, $p<0.01$). Apart from these differences in levels, there are few gender-specific differences in the multigenerational case. The patterns are largely similar for earnings ranks, with gender-specific transmission in the inter- but less so in the multigenerational case.141414engzell2023understanding and branden2024like show similar gender-specific patterns for intergenerational persistence in earnings ranks in Sweden. We study below if regional rankings are similar when considering different genders and lineages.
Table 2: Inter- and multigenerational correlations for different lineages
Unweighted Weighted
Yrs. education Earnings ranks Yrs. education Earnings ranks
Mean SD Mean SD Mean SD Mean SD
Father-child .270 .052 .214 .037 .298 .046 .223 .030
Mother-child .290 .044 .196 .036 .311 .038 .202 .028
Parental average-child .335 .048 .322 .044 .359 .041 .322 .034
Grandfather(pat.)-child .100 .052 .087 .040 .126 .054 .096 .030
Grandmother(pat.)-child .070 .044 .090 .039
Grandfather(mat.)-child .098 .047 .089 .037 .120 .044 .101 .031
Grandmother(mat.)-child .072 .038 .087 .035
Grandparental avg.-child .123 .051 .129 .064 .148 .048 .145 .053
Note: Means and standard deviations across municipalities for different inter- and multigenerational measures. We do not report earnings rank coefficients for grandmothers due to smaller samples. However, the grandparental average earnings rank includes all four grandparents (if they are above the lower earnings limit, see Section 2).
Second, we exploit the availability of three-generation data and compare the relative sizes of our inter- and multigenerational correlations. Recent studies on the national level find that multigenerational correlations tend to be larger than a simple iteration of parent-child correlations would suggest, i.e. intergenerational transmission cannot be approximated by a first-order Markov process; see AndersonSheppardMonden2017 and stuhler2024multigenerational for recent surveys. [^15] Moreover, see LindahlPalme2014_IGE4Generations, Dribe:2016aa, hallsten2017grand, Engzell:2020aa, Adermonetal2016, HelgertzDribe2021, ColladoOrtunoStuhlerKinship and hallsten2023shadow for multigenerational evidence from Sweden.
We find the same pattern on the regional level: in 77% of municipalities, the estimated grandfather-child correlation in schooling $\beta_{-2}$ is larger than the square of the father-child correlation $\beta_{-1}$,
$\Delta=\beta_{-2}-\beta_{-1}^{2}>0.$
This share increases to 88% when weighting each municipality’s gap $\Delta$ by the number of underlying observations (i.e., 88% of all individuals live in municipalities where $\Delta>0$). The gap $\Delta$ tends to be larger in more populous municipalities and in those exhibiting greater socioeconomic inequality in the parent generation (see Section 5). [^16] The gap $\Delta$ remains similar when using generation-specific averages to measure education, as this change increases $\beta_{-1}^{2}$ and $\beta_{-2}$ proportionally. The share with $\Delta>0$ is even larger when considering earnings rank coefficients, with 87% and 94% for the unweighted and weighted cases, respectively.
We further conduct formal statistical tests of $\Delta=0$ across municipalities. [^17] Since $\Delta$ is a non-linear function of parameters, we use the delta rule to derive its approximate standard error and t statistics separately for each municipality, $Var(\triangle)$ $\approx\left(\frac{\partial\triangle}{\partial\beta_{-2}}\right)Var(\beta_{-2})+\left(\frac{\partial\triangle}{\partial\beta_{-1}}\right)^{2}Var(\beta_{-1})+2\left(\frac{\partial\triangle}{\partial\beta_{-2}}\right)\left(\frac{\partial\triangle}{\partial\beta_{-1}}\right)Cov(\beta_{-2},\beta_{-1})$ $\approx Var(\beta_{-2})+(2\beta_{-1})^{2}Var(\beta_{-1})-2(2\beta_{-1})Cov(\beta_{-2},\beta_{-1})$ where we retrieve estimates of $Cov(\beta_{-2},\beta_{-1})$ from the robust covariance matrix generated by seemingly unrelated regressions (SUR). The test statistic is then given by $T=\frac{\triangle}{\sqrt{Var(\triangle)}}$ and can be compared to standard normal critical values. For education, the t statistics are larger than 1.96 in about 33% (unweighted) and 55% (weighted by the number of pairs) of municipalities. For earnings ranks, the corresponding shares are 42% and 67%. Figure A.3 in the Appendix plots the full distribution of the test statistics for both education and earnings ranks, showing that $\Delta$ is virtually never significantly negative. Consequently, the implied rejection rates for the one-sided hypothesis $\Delta\leq 0$ are high.
These results imply that parent-child correlations capture only part of the intergenerational process, and tend to understate the extent of status transmission in the “long run” (across many generations). But while understating the level of transmission, parent-child correlations may still capture differences in long-run transmission across regions. We test and confirm this hypothesis in Section 4. Conventional parent-child correlations therefore remain useful for comparative purposes, irrespectively of whether our interest centers on two- or multigenerational processes.
3.2 Mobility maps
Figure 2: Inter- and multigenerational education mobility across Swedish municipalities
(a) Father-child schooling correlation
(b) Grandfather-child schooling correlation
Notes: The figures plot estimates of the father-child correlation in years of schooling (sub-figure a) and grandfather-child correlation (sub-figure b), across 290 Swedish municipalities. The estimates are based on birth cohorts 1981-89.
To further illustrate how status transmission varies across municipalities, Figure 2 maps estimates of the father-child (sub-figure a) and the grandfather-child (sub-figure b) correlations in educational attainment. The point here is not to focus on any particular municipality, as the municipality-specific estimates can be quite noisy. Instead, we aim to show that both inter- and multigenerational mobility varies systematically between municipalities. The intergenerational correlation tends to be higher in more densely populated areas, and clusters of municipalities with high correlations (i.e. low mobility) are visible around Sweden’s major cities. [^18] The correlation between the father-child correlation in years of schooling and the number of father-child pairs observed in each municipality is 0.36.
In Figure A.4 in the Appendix we illustrate the corresponding geographical variation in earnings rank slopes. The spatial pattern of intergenerational mobility in earnings is somewhat different, but again tends to be higher in more populated municipalities in or around larger cities, though not to the same extent as for education. Figure A.5 in the Appendix shows the geographic distribution of two alternative measures of earnings mobility, the IGE and absolute upward (rank) mobility (for the intergenerational case only). The spatial pattern of upward mobility in earnings tends to be greater in some municipalities close to the larger cities, in central southern municipalities, and in some rural areas in northern Sweden.
We can compare our estimates to Heidrich:2017aa, who provides detailed national and regional estimates of intergenerational income mobility for the Swedish population born between 1968 and 1976. We consider more recent cohorts and our empirical specification differs in various aspects. In particular, we predict prime-age earnings based on all available earnings data of each individual, compared to using a conventional shorter-run average of observed earnings or income. Another important difference is that we consider variation on the finer municipality level, while Heidrich focuses on the broader definition of local labor market regions. [^19] One could debate which level of aggregation makes more sense. Child care, compulsory school and high schools are organized at the municipality level, and financed by municipal income taxes. Many other local amenities and services are also best distinguished at the municipal level. Local labor market regions are larger and defined based on commuting patterns. Different divisions are thus likely to capture differences in mechanisms underlying intergenerational transmission to different extents. The exact division is not key in our setting, as we primarily aim at making a set of conceptual and methodological points. Consistent with this and other differences in specification, we find slightly higher earnings persistence. Nevertheless, the spatial pattern appears largely similar. For example, Heidrich:2017aa notes that several local labor markets in the south of Sweden and around Stockholm are doing particularly well in lifting children from lower income families, while absolute upward mobility is particularly low in municipalities in the Swedish inland located close to the Norwegian border. These patterns are also visible in our Appendix Figure A.5b. Given that we study different birth cohorts, this similarity also implies that mobility differences between regions are systematic and at least somewhat persistent, rather than just temporary.
3.3 Linear approximations
Regional comparisons of intergenerational mobility, such as those reported in Figures 1 and 2, are subject to two methodological concerns. The first is whether intergenerational relations can be sufficiently well summarized by simple summary measures, such as linear regression or correlation coefficients. CholliDurlauf2022 note that theoretical models usually imply nonlinear relationships between parent and child status, which are often ignored in practice.
For illustration, Figure 3a plots the mean child earnings rank against the parental earnings rank at the national level. The rank-rank relation is linear over most of the distribution, but much steeper at the very bottom and top. [^20] NybomStuhlerJHR2017 document a similarly steep relation at the very bottom of the parental income distribution. The conditional expectation function (CEF) is therefore not fully captured by a linear regression, in contrast to the pattern for the U.S. as described by CHKS_Where_QJE. However, it serves as a good approximation over most of the distribution, such as between the 10th and 90th percentiles. The expectation of child years of schooling conditional on father’s years of schooling is approximately linear for most of the distribution as well (see Appendix Figure A.6).
Figure 3: Non-linear intergenerational earnings rank slopes
(a) National level
(b) Separately for the five largest cities
Notes: Binned scatter plot of the expectation of child earnings rank conditional on father’s earnings rank on the national level (sub-figure a) or municipal level (sub-figure b). Ranks are defined relative to the national distribution.
While the CEF appears sufficiently linear on the national level, it does not necessarily follow that the conditional expectations are equally linear on the regional level. For illustration, Figure 3b plots the mean child earnings rank conditional on parent rank deciles for Sweden’s five largest municipalities. Upward mobility is substantially lower, and the rank-rank slope steeper, in Malmö (diamonds) compared to for example Uppsala (grey circles). However, apart from such differences for the bottom decile, the relationships are approximately linear in all cities; while it is still somewhat steeper at the bottom and the top of the parental earnings distribution, this pattern is less pronounced at the regional level. We provide the corresponding graphs for the grandfather-child relationship in Figure A.7 in the Appendix. The region-specific multigenerational CEFs are approximately linear, and less steep at the bottom of the income distribution than their national-level counterpart. [^21] Further, we conducted a simple analysis for all municipalities by comparing the regression $R^{2}$ from municipality-specific linear and quadratic regressions. The quadratic model generates an $R^{2}$ that on average is only 2.7% larger than the mean $R^{2}$ from the linear model. This number drops to only 1.4% if we weight the averages by number of individuals rather than by municipality. Linear measures therefore appear well suited for cross-regional comparisons.
3.4 Sampling error
A second concern in fine-grained regional comparisons is sample size: Some of the variation across municipalities might not be “real” variation but reflect sampling error, in particular for smaller municipalities for which comparatively few parent-child pairs are observed. [^22] As we observe (almost) the entire population, our estimates do not suffer from traditional sampling error. However, the observed mobility for a given set of parent-child pairs can be interpreted as a random draw from the “potential” distribution for that municipality. While the mean number of father-child pairs per municipality for children born between 1981 and 1989 is 3,183 for education and 3,099 for earnings, in 57 of the municipalities (59 for earnings) we observe fewer than 1,000 pairs. [^23] Moreover, inference based on earnings ranks is complicated by the fact that each unit’s rank depends on which other units have been sampled (MogstadWilhelm23). We also observe fewer grandfather-child pairs than father-child pairs, and especially so for earnings. Education is observed for all who have resided in Sweden at some point as adults, while earnings is only observed for those who at some point from 1968 have recorded taxable labor earnings above a certain lower amount (see Section 2). To probe the influence of sampling error, we implement two tests.
Figure 4: Intergenerational correlations by municipality size
Notes: Density plot of the estimated intergenerational correlations in years of schooling for father-child pairs across 160 “small” municipalities ($\leq$ 2,000 observations), 130 “large” municipalities (>> 2,000 observations), and corresponding “placebo” municipalities in which we randomly allocated pairs across municipalities.
First, we split the sample into smaller ($\leq$ 2,000 observations) and larger municipalities (>> 2,000 observations), with the expectation that the influence of sampling error should be more limited for the latter. Figure 4 plots the resulting densities of the father-child correlation in years of schooling. The correlations vary nearly as much for large as for small municipalities (the standard deviations are 0.038 vs. 0.048). Figure A.8 in the Appendix shows corresponding distributions for earnings. [^24] Figure 4 shows that the educational correlations are on average substantially lower for small municipalities, a pattern that is possibly related to the greater homogeneity in educational attainment in smaller municipalities. And while the estimated coefficient in an OLS regression of the municipal parent-child correlation on sample size is large and significant, it loses significance if we control for the standard deviation of years of schooling. The difference in means between small and large municipalities for earnings ranks is smaller, but qualitatively similar (see Figure A.8 in the Appendix). Second, we compute “placebo correlations” by reshuffling parent-child pairs randomly across all 290 municipalities, such that variation in the placebo correlations solely reflects chance. Figure 4 shows that for both small and large municipalities, the resulting density is much narrower than the actual variation of intergenerational correlations across municipalities. Moreover, there is considerably less dispersion in the placebo distribution for large than for small municipalities – the standard deviation of the former is less than half of the one for the latter. [^25] Considering instead changes, the decrease in dispersion is larger for large than for small municipalities when comparing the placebo with the true distributions, with a drop in the std. dev. of about 70% compared to 45% for small municipalities. These differences between small and large municipalities are even more pronounced for earnings ranks. In particular, the std. dev. of the placebo distribution of earnings rank slopes of large municipalities is about 44% of the one for small ones.
Still, some of the reported variation between municipalities reflects noise, in particular for small municipalities, and sampling error might explain much of the regional variation in studies that use less data than we do here. Sampling error may have a greater effect for earnings-based measures, as earnings are themselves subject to transitory noise (mazumder2005fortunate). This is the case in our setting as well, even though we use comparatively high-quality earnings measures (see Appendix Figure A.8). Few studies have addressed this issue systematically. Rather than estimating separate regressions for each region, Heidrich:2017aa estimates a multilevel model that explicitly accounts for the influence of sampling error on the variation of regional mobility estimates. This approach effectively “shrinks” the region-specific estimates to their mean, in particular for small regions with imprecise estimates. Risa:2019aa implements a similar shrinkage procedure for mobility estimates for Norway. In addition, Heidrich:2017aa reports only those estimates that are significantly differently from the mean, replacing all other region-specific estimates with that mean.
In this paper, we instead focus on the raw region-specific estimates, as our aim is to compare different mobility measures (Section 4) and to relate our findings to the existing literature reporting raw correlations. Moreover, we will relate those region-specific mobility estimates to other regional characteristics, such as inequality (Section 5). To reduce the influence of sampling error, we weight all regressions by the number of underlying parent-child pairs for each region, and restrict parts of the analysis to larger municipalities with more precisely estimated statistics. However, we also report unweighted estimates for our core analyses. As illustrated in Appendix Figure A.1, a substantial share of individuals live outside of the largest municipalities. Thus, while our weighted estimates give more influence to larger municipalities, they are not dominated by them.
4 Are Regional Rankings Stable?
Many different types of inter- and multigenerational mobility statistics are in use, but the researcher’s choice of which statistic to report is typically restricted by data availability. Intergenerational linked income data is not available for many countries, and most data sets contain data on only two generations. It is therefore important to understand whether regional comparisons are sensitive to the choice of mobility statistic on which they are based. Specifically, we study whether regional comparisons are sensitive to the choice of outcome (education vs. earnings), time frame (inter- vs. multigenerational), or lineage (paternal vs. maternal line).
Our main focus is on whether intergenerational (parent-child) associations – which can be estimated in many settings – reflect similar underlying mobility processes as more data-demanding multigenerational associations. This analysis is complementary to a recent analysis by deutscher2023measuring, who compare the ranking of Australian regions across different measures of intergenerational income mobility, including measures of relative and absolute mobility, and more complex measures of “inequality of opportunity”. Our main contribution is to compare rankings across parent-child and grandparent-child correlations. We present our analyses for three statistics, correlations in education and slope coefficients in earnings ranks and log earnings (the IGE).
Figure 5: Different measures of intergenerational mobility across regions
Notes: Scatter plots of different intergenerational dependence measures across Swedish municipalities. We compare the father-child correlation in years of schooling with the paternal grandfather-child correlation (top-left panel), the mother-child with the maternal grandfather-child correlation in schooling (top right), and similarly for slope coefficients in earnings ranks (center left and center right) and log earnings (bottom left and bottom right). The plots are restricted to municipalities with at least 1,000 parent-child pairs with observed schooling.
Figure 5 illustrates our findings with scatterplots including labels for the five most populous municipalities. We complement the figure with a full correlation matrix, shown in Table B.1 in the Appendix. The top panel of Figure 5 illustrates that regions rank similarly irrespectively of whether we consider the intergenerational correlation in years of schooling (x-axis) or the corresponding multigenerational correlation (y-axis). The pattern is similar along both the paternal and maternal line, though the latter appears slightly more dispersed. The correlation between the two measures weighted by sample size is 0.79 along the paternal and 0.72 along the maternal line. The pattern reflects that mother-child and father-child correlations are themselves strongly related (weighted correlation between the two at 0.83). This finding is intuitive in the Swedish context, but may not extrapolate to other settings or other outcomes in which gender differences are more pronounced.
The middle panels show that municipalities also rank similarly across inter- and multigenerational correlations when considering earnings ranks, even though these patterns are somewhat more dispersed. The weighted correlation for the paternal line (mid-left panel) is 0.51, while it is 0.48 along the maternal line (mothers and maternal grandfathers). Finally, the bottom panel shows the same plots for the IGE based on log earnings. The patterns are similar to those found for earnings ranks, though the implied correlations are somewhat weaker (at 0.47 and 0.43, respectively). These observations suggest that conventional parent-child correlations remain a useful summary measure for mobility differences across regions, even though they understate the overall transmission of advantages in the long run (Section 3). Still, a comparison of inter- and multigenerational correlations may provide additional insights into why mobility differs, as we show in Section 6.
Moreover, Figure 5 indicates that educational and earnings mobility are not strongly related; compare for example the five largest municipalities that are labeled in the graph in the upper with the lower two panels. Detailed correlations are shown in Table B.1 in the Appendix. Municipalities that have low parent-child educational mobility tend to also have lower mobility in earnings, but this relation is weak (the correlation is 0.06 for fathers and 0.18 for mothers). The relation tends to be even weaker, and in some cases negative, when considering the intergenerational elasticity of earnings (IGE) rather than earnings ranks. This observation stands in contrast to cross-national comparisons, in which educational and income mobility are more strongly related (Blanden2011Survey, Stuhler2018JRC). One likely reason is that regional measures of income mobility can be quite noisy (Heidrich:2017aa). Indeed, we find that the correlation between educational and income mobility becomes more sizable when restricting our sample to municipalities with larger sample sizes. And interestingly, the relation between education and earnings mobility is generally higher for multi- compared to intergenerational measures (correlations around 0.3 along both the paternal and maternal lines).
Finally, we can compare different measures of earnings mobility. Municipalities characterized by high rank correlations in earnings also have a high IGE. As shown in Table B.1 in the Appendix, this relationship between different earnings measures is similarly strong for inter- and multigenerational correlations. Some of our analyses here are similar to engzell2023understanding, who study the relationships across a multitude of different measures of intergenerational mobility in Sweden using variation over cohorts.
We conclude that inter- and multigenerational mobility measures are closely related, with regions characterized by high parent-child mobility also demonstrating high three-generation mobility. In contrast, education- and earnings-based mobility measures are generally not as close substitutes. Consistent with this conclusion, we observe that the largest Swedish cities exhibit opposing patterns in educational and earnings mobility. While Malmö (the third largest city) is characterized by low earnings mobility, it ranks closer to the Swedish average in terms of educational mobility. Uppsala (the fourth largest city) instead is characterized by very low educational mobility and higher than average earnings mobility. As we saw in Figure 3b, Malmö is an outlier also in terms of absolute upward mobility. The expected earnings rank of children at the 25th percentile of the paternal distribution is 0.43, which places the city in the bottom decile of all municipalities. [^26] An open question that we do not explore further here is whether observed earnings for residents in Malmö might be affected by cross-border commuting to Copenhagen. Still, this rate of upward mobility compares favorably to many regions in the US, in which the expected earnings rank at the 25th percentile can be as low as 0.36 (CHKS_Where_QJE).
5 The Great Gatsby Curve Across Regions
How intergenerational mobility varies across regions is interesting from a purely descriptive perspective. However, such regional variations might also point to specific correlates or causal determinants of mobility, providing insight into how regions with high or low intergenerational mobility differ from other regions.
We focus here on the relation between inter- and multigenerational mobility with cross-sectional inequality. A robust pattern in international comparisons is that intergenerational income mobility and inequality are negatively associated across countries (Blanden2011Survey; CorakJEP2013). In addition, many theories on intergenerational transmission predict that income persistence varies across the income distribution (durlauf2022great). One particular mechanism contributing to both intergenerational and cross-sectional inequality is the segregation of families into economically homogeneous neighborhoods (Durlauf1996).
However, it is unclear whether this statistical relation between inequality and mobility, sometimes labelled as the “Great Gatsby Curve”, reflects a direct causal link, a third factor affecting both inequality and mobility, or whether the association is entirely spurious. Regional comparisons are a useful next step, as the number of observations can be much larger than in cross-country comparisons, as mobility statistics are from the same data source and therefore more comparable, and as many institutional and economic factors are held fixed when comparing regions within the same country. We here extend this regional approach by analyzing the relation between inequality and mobility across multiple generations.
Figure 6: Great Gatsby Curves across Swedish municipalities
(a) Intergenerational Gatsby Curve
(b) Multigenerational Gatsby Curve
Notes: The figure plots (subfigure a) the intergenerational rank slope in earnings against a measure of earnings inequality among fathers (the Gini index) and (subfigure b) the multigenerational (child vs. paternal grandfather) rank slope in earnings against earnings inequality among the grandfathers. Scatter plots and linear fit restricted to municipalities for which at least 1,000 father-child pairs are observed. The linear regression line is weighted by the number of father-child pairs.
Figure 6 plots the inter- (father-child, subfigure a) and multigenerational (grandfather-child, subfigure b) rank slopes in earnings against a measure of earnings inequality. For the intergenerational analysis, we measure inequality by the Gini index among fathers in the parent generation, while in the three-generation case we use the Gini index among grandfathers. The results are similar when considering alternative inequality measures, such as the standard deviation of earnings. We find a weak but positive relation, with those municipalities characterized by high cross-sectional inequality being on average more likely to be characterized by low intergenerational mobility as well.
Figure 6 also documents a multigenerational Gatsby curve. Those municipalities that had more inequality in the grandparent generation also tend to have lower grandchild-grandparent earnings (rank) mobility. While the relationship is fairly weak, it is about as strong as in the intergenerational case. Further, the association between multigenerational mobility and inequality becomes much stronger if we measure inequality in the parent rather than the grandparent generation (see Appendix Figure B.1). While there is no apriori reason to prefer the former over the latter way to measure inequality, we focus on grandparental inequality in our baseline such that inequality is measured before mobility.
Table 3: The Great Gatsby Curve across Swedish municipalities
Sons Daughters Sons & Daughters
(1) (2) (3) (4)
Control for municipality size x
(A) Pairwise correlations between intergenerational measures and the Gini
Earnings
Rank slope $0.086$ $0.191$ $0.172$ $0.235$
$p=0.143$ $p=0.001$ $p=0.003$ $p=0.001$
IGE (log) 0.074-0.074 0.052-0.052 0.078-0.078 $0.012$
$p=0.207$ $p=0.373$ $p=0.184$ $p=0.844$
P25 upward rank mobility 0.356-0.356 $0.207$ 0.125-0.125 0.048-0.048
$p=0.000$ $p=0.000$ $p=0.033$ $p=0.412$
Education (yrs.)
Correlation $0.580$ $0.495$ $0.584$ $0.632$
$p=0.000$ $p=0.000$ $p=0.000$ $p=0.000$
(B) Pairwise correlations between multigenerational measures and the Gini
Earnings
Rank slope $0.077$ $0.181$ $0.153$ $0.145$
$p=0.194$ $p=0.002$ $p=0.009$ $p=0.013$
IGE (log) 0.024-0.024 $0.033$ $0.001$ $0.012$
$p=0.682$ $p=0.576$ $p=0.991$ $p=0.840$
P25 upward rank mobility 0.117-0.117 $0.351$ $0.097$ $0.285$
$p=0.046$ $p=0.000$ $p=0.101$ $p=0.000$
Education (yrs.)
Correlation $0.684$ $0.603$ $0.717$ $0.697$
$p=0.000$ $p=0.000$ $p=0.000$ $p=0.000$
Notes: The table reports the pairwise correlations between the indicated inter- or multigenerational statistic and the Gini index (among fathers in Panel A, among grandfathers in Panel B), weighted by the number of father-child pairs. The rank slope is the slope coefficient from a regression of child on father’s (Panel A) or grandfather’s (Panel B) earnings rank. The IGE (log) is the slope coefficient from a regression of child log earnings on father’s (Panel A) or grandfather’s (Panel B) log earnings. P25 upward mobility is defined as the expected child earnings rank at the 25th percentile of father’s (Panel A) or grandfather’s (Panel B) ranks. Columns (1)-(3) report raw correlations for sons, daughters, and sons and daughters pooled, while column (4) reports the correlation between the residuals from separate regressions of the Gini and intergenerational coefficient on sample size.
How robust is this association between earnings inequality and mobility to the choice of mobility measure and generational perspective (see Section 4)? Table 3 extends on Figure 6 by reporting the correlation between different measures of inter- and multigenerational mobility and the Gini index. We report estimates separately for subsamples of father-son and father-daughter pairs (Columns (1) and (2)) and a pooled sample comprising both sons and daughters (Columns (3) and (4)). For the latter, the first column reports raw correlations while the second column controls for municipality size, as income inequality tends to be more pronounced in the larger cities. [^27] Specifically, we regress the Gini index and each inter- or multigenerational statistic on the number of complete father-child pairs and report the pairwise correlations between the residuals from those regressions. All correlations are weighted by the number of parent-child pairs observed for each municipality.
The first row of Table 3 reports the full-sample equivalent to subfigure a in Figure 6, correlating the rank slope in earnings with the Gini coefficient across all 290 Swedish municipalities. The correlations are close to 0.2 when considering the pooled sample and father-daughter pairs, but lower and insignificant for father-son pairs. [^28] Given the that correlations are constrained within [−1,1] we report p-values rather than standard errors; these p-values do not account for the fact that the underlying statistics are themselves estimated. The correlations are close to zero and insignificant when considering the intergenerational elasticity of earnings (i.e., the log-log regression slope), which might partly stem from that the IGE is mechanically negatively (positively) related to the variance of father’s (child’s) log earnings. [^29] Another approach would be to use the mean of the Gini indices in the father’s and the child’s generations, and thus isolate the analysis from this mechanical relationship (as in e.g. CorakJEP2013). One could also standardize log earnings within generations and relate the intergenerational correlation in log earnings with the Gini index. In our data, both these approaches result in positive and sometimes statistically significant relationships (not shown here). The third row shows that, among sons as well as in the pooled sample, the expected earnings rank of those at the 25th percentile of the parental distribution is significantly lower when income inequality is high. Surprisingly, this relation switches sign when only considering daughters, presumably due to female labor force participation being higher in more unequal municipalities. However, apart from this latter result and the results based on the IGE, we tend to confirm the pattern documented in the discourse on the Great Gatsby Curve: intergenerational persistence tends to be systematically higher, and upward mobility significantly lower (for sons), when earnings inequality is high.
This negative relation between income inequality and intergenerational mobility also holds when measuring mobility based on educational outcomes, and indeed is much more pronounced. For example, the cross-regional correlation between the Gini index and the father-child correlation in years of schooling is around 0.5-0.6, depending on specification. The intergenerational evidence in Panel A is generally also robust to controlling for municipality size (column 4), suggesting that the driver of the mobility-inequality relationship is not simply due to larger municipalities being more unequal.
We next consider the Gatsby curve over three generations, which is the more novel feature of our analysis. As shown in Panel B of Table 3, the Gini index (now defined among the grandfathers) relates also to the multigenerational grandfather-child correlations, following a largely similar pattern as in the intergenerational case. There is again a negative relationship between inequality and mobility in terms of both earnings ranks and schooling, and the relationship with absolute upward mobility diverges completely between sons and daughters. The relationships tend to be somewhat weaker for earnings ranks but slightly stronger for educational mobility, compared to the intergenerational case. While the table reports only the parent- and grandparent-child correlations from the paternal side, the pattern is similar for other possible gender combinations.
This evidence adds to a small number of recent studies that document a similar “Great Gatsby Curve” across regions within countries, although rarely for the three-generation case. [^30] Examples include CHKS_Where_QJE for the U.S. or GuellPellizzariPicaMora2018 for Italy, among many others. For Sweden, Branden2019 shows that men who grew up in regions with high levels of income inequality experienced less intergenerational mobility, and highlights the role of educational attainment and cognitive and non-cognitive skills in this relationship. Heidrich:2017aa also provides evidence on a “Great Gatsby Curve” across regions within Sweden. We find that the strength of within-country “Great Gatsby Curves” depends on the choice of mobility statistic and how and in what generation inequality is measured. Such choices can also substantially impact the magnitudes of the Gatsby relationship. We further find that the link between intergenerational mobility and income inequality is even more pronounced for educational outcomes than for earnings, which is notable given that education and earnings mobility statistics are only imperfectly related (see Section 4). These patterns are replicated, and the association between income and educational mobility even stronger, in the multigenerational case.
6 Interpreting Regional Differences in Mobility
Why does intergenerational mobility vary across regions? Following CHKS_Where_QJE, the literature has searched for regional correlates that can explain some of the regional variation in mobility. This literature has been fruitful, in that it has identified various factors associated with low mobility, which can then be studied further in more targeted, causal research designs.
In this section, we follow a related but different strategy, which draws on insights from the recent multigenerational literature. Rather than searching for specific observable characteristics to explain mobility differentials, we compare inter- and multigenerational correlations to gain a deeper insight into how and why mobility varies across municipalities. In particular, we will be able to distinguish whether regional differences in mobility are “short-term” phenomena affecting mobility between parents and children, or whether they are even more pronounced in terms of long-run mobility across many generations. While we cannot directly map our estimates into any particular mechanism, our approach provides new “statistical facts” about mobility differentials across regions that indirectly are informative about which types of mechanisms can or cannot explain them.
To provide a simple example, imagine that income mobility might vary across regions either because the returns to human capital differ (e.g., because of different industrial composition) or because the intergenerational transmission of human capital differs (e.g., because of differences in residential segregation and the quality of the schools and teachers; see e.g. Durlauf1996). In a simple theoretical model of transmission, the latter would have more persistent effects on the prospects of families with low human capital (as we show below). By comparing regional differences in inter- and multigenerational correlations researchers might be able to distinguish these two stories.
6.1 A simple latent factor model
We formalize these ideas in a simple latent factor model. Assume transmission in generation $t$ of family $i$ in region (municipality) $r$ is governed by
$y_{it}$ $=\rho_{r}e_{it}+u_{it}$ (2)
$e_{it}$ $=\lambda_{r}e_{it-1}+v_{it},$ (3)
where observed outcome $y$ depends on latent “endowments” $e$ according to returns $\rho_{r}$, which are partially transmitted within families according to transferability $\lambda_{r}$, and where $u$ and $v$ are white-noise error terms representing market and endowment luck, uncorrelated with each other and past values.
To simplify the presentation we drop the i subscript and assume that e and y are standardized within each region with mean zero and variance one, such that the slopes $\rho_{r}$ and $\lambda_{r}$ can be interpreted as correlations. For example, if the outcome of interest is earnings, the parameter $\rho_{r}$ would capture the fraction of earnings that in region $r$ is explained by an individual’s own human capital or characteristics, as opposed to factors or events outside of the individual’s control, such as market luck or market-level shocks; and $\rho_{r}=1$ would imply that earnings differences are fully explained by an individuals’ own characteristics. [^31] As already mentioned, it can be difficult to map specific mechanisms into the parameters of this model; the approach is therefore a complement, not a substitute for the popular approach of searching for specific “correlates” of mobility across regions.
What are the implications of this model for inter- and multigenerational correlations? Given this model, the parent-child correlation in $y$ in region $r$ equals
$\beta_{-1,r}$ $=\frac{Cov(y_{t},y_{t-1})}{Var(y_{t-1})}=\frac{Cov(\rho_{r}e_{t}+u_{t},\rho_{r}e_{t-1}+u_{t-1})}{1}$
$=\rho_{r}^{2}Cov(\lambda_{r}e_{it-1}+v_{it},e_{t-1})=\rho_{r}^{2}\lambda_{r},$ (4)
illustrating that regional variations in intergenerational mobility – as documented for an increasing set of countries – could stem from variation in “returns” $\rho_{r}$ or variation in “transferability” $\lambda_{r}$. [^32] As noted by a referee, substituting eq. (3) into (2) and combining the resulting expression with $\lambda_{r}$ times the lagged value of (2) yields the dynamic equation $y_{it}=\lambda_{r}y_{it-1}+\eta_{it}$, where $\eta_{it}=\rho_{r}v_{it}+u_{it}-\lambda_{r}u_{it-1}$. Applying the omitted variable formula to account for the correlation between $y_{it-1}$ and $u_{it-1}$, we once again derive solution (4). Based on intergenerational correlations alone, one cannot determine which of these factors is more important. However, following the same solution steps, the correlation across three generations in this model equals
$\beta_{-2,r}$ $=\frac{Cov(y_{t},y_{t-2})}{Var(y_{t-2})}=\rho_{r}^{2}\lambda_{r}^{2}$ (5)
illustrating that any further decay of status advantages over more than two generations depends only on the “transferability” of endowments $\lambda_{r}$ and not on their returns $\rho_{r}$. In other words, different components of the transmission process affect inter- and multigenerational correlations differently, and their comparison therefore tells us something new about the source of mobility differences across regions. [^33] Our argument here is an example for the more general insight that the observation of the dynamics of an outcome variable can be helpful in parameter identification of its underlying model (AignerHsiaoKapteynWansbeek1984).
Formally, from equations (4) and (5) it follows
$\frac{\beta_{-2,r}}{\beta_{-1,r}}=\lambda_{r}$ (6)
and
$\left(\frac{\begin{array}[]{c}\beta_{-1,r}\end{array}^{2}}{\beta_{-2,r}}\right)^{1/2}=\rho_{r}$ (7)
so that region-specific estimates of $\lambda_{r}$ and $\rho_{r}$ can be constructed from our estimates of inter- and multigenerational correlations as shown in Section 3.
Table 4: Regional estimates of $\lambda_{r}$ and $\rho_{r}$
Education Earnings
$\hat{\rho}_{r}$ $\hat{\lambda}_{r}$ $\hat{\rho}_{r}$ $\hat{\lambda}_{r}$
Panel A: Baseline sample
Mean 0.890 0.403 0.756 0.431
Std. Dev. 0.244 0.122 0.574 0.131
Panel B: Balanced sample
Mean 0.924 0.391 0.731 0.442
Std. Dev. 0.252 0.118 0.238 0.133
Notes: The table reports the mean and standard deviation of estimates of the “returns” $\rho_{r}$ and the “transferability” $\lambda_{r}$ across 290 municipalities, weighted by grandfather-child pairs in each municipality.
A key property of the latent factor model is that it can explain why the estimated grandfather-child correlation in schooling $\beta_{-2}$ is larger than the square of the father-child correlation $\beta_{-1}$,
$\Delta=\beta_{-2}-\beta_{-1}^{2}>0,$
as is the case for most Swedish regions (see Section 3). As this finding of “excess persistence” $\Delta>0$ is common, the model has been used in several recent multigenerational studies (e.g., ECOJ:ECOJ12453; NeidhoeferStockhausen; Bellocetal2024; Celhaygallegos2025).
Of course, this model is highly stylized, abstracting from many mechanisms that one might want to study explicitly. Its one-parent structure abstracts from the role of assortative mating, although the parameter $\lambda_{r}$ can be interpreted as a reduced-form representation of the combined effects of intergenerational transmission and assortative mating (ECOJ:ECOJ12453). It also ignores that the intergenerational transmission process may differ by gender (e.g., pembrey2006sex) or neighborhood or peer effects. Perhaps most importantly, we assumed that the underlying distributional moments are in a steady state, which is implausible for data spanning three generations (especially given the large shifts in means shown in Table 1). In our application, we abstract from shifts in the variance across generations by focusing on correlation rather than regression coefficients. Moreover, we define $\beta_{-1}$ as an appropriately weighted average of the correlations between children and parents (G1-G2) and between parents and grandparents (G2-G3). [^34] Specifically, $\beta_{-1}=\sqrt{\beta_{-1}^{G1-G2}\beta_{-1}^{G2-G3}}$. See AtkinsonJenkins1984 and NybomStuhler2019 for a discussion of steady-state assumptions in mobility research. ECOJ:ECOJ12453 discuss the estimation of a latent factor model with time-varying parameters.
Still, the model captures the notion that an individual’s observed outcome $y$ might be a poor proxy for other endowments and advantages that influence the prospects of one’s child; and it shows that this simple notion has important implications for the relative size of inter- and multigenerational correlations. Extended versions of this model can explain the pattern of intergenerational transmission across different kins (ColladoOrtunoStuhlerKinship) or rationalize why intergenerational transmission is higher at the surname than the family level (Clark2014book).
One obvious extension would be to allow for direct effects of parental outcomes $y_{t-1}$ on child outcomes $y_{t}$ in eq. (2), as in becker1986human or solon2004model. However, a commonly used version of this model as described in solon2004model implies that multigenerational correlations should be smaller than one might expect from parent-child correlations, i.e. $\Delta<0$, and is therefore less consistent with the observed multigenerational patterns. [^35] Put differently, this model implies that in a regression of child on parent and grandparent status, the coefficient on grandparents is negative; see Solon201413 for details. And while studying more general versions of the Becker-Tomes model would be interesting, we do not pursue this path here, as with three generations there is not sufficient empirical information to identify the parameters of such extended models (see ColladoOrtunoStuhlerKinship, for a more general model fitted on more extensive set of kinship moments).
We estimate the latent factor model using either education or earnings as our variable of interest. In Table 4, we describe the resulting estimates for these two outcomes, and in two different samples. In Panel A we consider all father-child and grandfather-child pairs while in Panel B we consider a balanced sample, in which we only retain those observations for which we observe complete child-father-grandfather triplets. For simplicity, and to in line with the one-parent structure of the model, we consider father and paternal grandfathers in this estimation, but the results are similar when considering maternal lines.
We find that on average, estimates of the “returns” $\rho_{r}$ are around 0.9 for education and about 0.75 for earnings. Estimates of the “transferability” $\lambda_{r}$ are much smaller, around 0.4 for both outcomes. This pattern is very similar in our baseline and balanced samples. We further find that estimates of both $\rho_{r}$ and $\lambda_{r}$ vary substantially across municipalities. Of course, some of this variation will be due to sampling error, as already noted in Section 3.4. We will return to this issue below.
6.2 Regional differences in intergenerational mobility in the latent factor model
We next interpret regional differences in intergenerational mobility, as documented in Section 3, through the lens of the latent factor model. Table 5 shows estimates from a regression of the grandfather-child correlation in years of schooling (as plotted in the center-left panel of Figure 5) on our education-based measures of returns $\rho_{r}$ and transferability $\lambda_{r}$ (as described in the top-left panel of Table 4). [^36] While we report robust standard errors here, cluster-robust standard errors on the level of 25 aggregate regions (defined by the first two digits of the municipality codes) are similar.
Table 5: Regional variation in mobility and the latent factor model
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$
Years of schooling Earnings ranks
(1) (2) (3) (4) (5) (6)
$\hat{\rho}_{r}$ (“returns”) -0.089*** 0.033*** -0.014* 0.001
(0.024) (0.013) (0.006) (0.002)
$\hat{\lambda}_{r}$ (“transferability”) 0.342*** 0.418*** 0.203*** 0.206***
(0.014) (0.020) (0.009) (0.012)
# Municipalities 280 290 280 284 289 283
Adj. $R^{2}$ 0.250 0.855 0.883 0.069 0.745 0.719
Notes: The dependent variable is the grandfather-child correlation in years of schooling (columns 1-3) or the grandfather-child correlation in earnings ranks (columns 4-6) across 290 Swedish municipalities. The independent variables are estimates based on the corresponding parent-child and grandparent-child correlations, see equations (6) and (7). Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.001.
We find that the estimates of the returns $\hat{\rho}$ are negatively correlated with the grandparent-child correlation, both when considering educational outcomes (column 1) or earnings (column 4), with $R^{2}s$ below 0.3. In contrast, estimates of the transferability $\hat{\lambda}$ are positively correlated with the grandparent-child correlation (columns 2 and 5), and explain most of the regional variation in mobility ($R^{2}>0.7$). In interpreting these results we need to take into account that the estimates $\hat{\rho}$ and $\hat{\lambda}$ are negatively correlated across municipalities. [^37] It is straightforward to show that sampling error would introduce such negative correlation, i.e. the negative correlation might not capture the actual relationship between $\lambda_{r}$ and $\rho_{r}$ across regions. In particular, this explains why the coefficient estimates in columns 1 and 4 are negative, even though $\rho$ in itself has of course a positive effect on the grandparent-child correlation (see equation (5)). When including both $\hat{\rho}$ and $\hat{\lambda}$ jointly, we find positive coefficients; again, $\hat{\lambda}$ has a stronger association with the grandfather-child correlation.
The regional variation in intergenerational mobility is therefore primarily due to variation in the transferability $\lambda$, not due to variation in returns $\rho$. This implies that the regional differences in mobility are really due to differences in the transmission process itself, rather than variation in how latent endowments map onto observed advantages in the data. And this in turn tells us that regional differences in parent-child correlations do in fact understate differences in long-run mobility across regions. Put differently, regional variations in mobility appears to primarily reflect “real” differences in the intergenerational transmission from parents to children as opposed to “superficial” differences in the relation between observed outcomes and latent advantages.
Of course, it is difficult to map these insights from our stylized model into specific causal mechanisms. They provide some indications; for example, based on our results it seems unlikely that differences in earnings mobility across regions would primarily reflect spatial variation in returns to schooling. Still, our results are compatible with different interpretations. For example, we cannot distinguish if regional differences in the transferability $\lambda$ are persistent attributes of those regions, or if they reflect a temporary influence of certain economic events or policies. Despite these limitations, we hope our findings illustrate the potential value of measuring multigenerational correlations in regional comparisons or similar comparative designs.
6.3 Regional differences in cross-sectional inequality in the latent factor model
Table 6: Regional variation in cross-sectional inequality and the latent factor model
Dependent variable: Gini index in earnings
(1) (2) (3)
$\hat{\rho}_{r}$ (“returns”) -0.016*** 0.012
(0.004) (0.009)
$\hat{\lambda}_{r}$ (“transferability”) 0.067*** 0.093***
(0.013) (0.021)
# Municipalities 280 290 280
Adj. $R^{2}$ 0.049 0.240 0.259
Notes: The dependent variable is the Gini index among fathers in the parent generation. The independent variables are estimates based on the father-child and grandfather-child correlations in years of schooling, see equations (6) and (7). Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.001.
We can perform a similar analysis to understand whether regional variations in cross-sectional inequality are due to differences in the transferability $\lambda_{r}$ of endowments from parents to children or differences in the returns $\rho_{r}$ to those endowments. Table 6 shows that, similarly as intergenerational measures, differences in inequality across municipalities are more affected by variation in the transferability of endowments (cf. regression R2 in columns 1 and 2, or the size of the coefficients in column 3). This reinforces the idea that cross-sectional and intergenerational measures are closely intertwined, as already explored in Section 5. Moreover, this may suggest that it is really differences in intergenerational transmission between parents and children that are underlying this relationship. Our findings here also explain why multigenerational measures are as strongly (or more strongly in the case of education) associated with inequality as parent-child measures (see Table 3): while parent-child correlations are larger, multigenerational correlations provide a more concentrated signal of differences in $\lambda_{r}$ (cf. equation (4) and (5)).
6.4 Robustness and limitations
[…]
Two important limitations in our analysis are the limited number of observations available per region, and that we allocate each family (lineage) to the children’s municipality at age 16, even if the parent or grandparent generation lived in another municipality. We consider these issues in more detail below. A third important limitation is that our earnings measures are unlikely to be fully comparable across generations. [^39] Table A.1 in the Appendix compares statistics based on our preferred measure of predicted earnings at age 40 with more standard measures based on short-run averages. While most combinations of measures yield largely similar dispersion across areas, our preferred measure gives larger coefficients that are more closely aligned with the recent mobility evidence for Sweden.
Sampling error. One important caveat in these results is that they also reflect the influence of sampling error. Our estimates of inter- and multigenerational correlations contain noise (as shown in Section 3.4), and this will map into noise in our estimates of $\rho_{r}$ and $\lambda_{r}$. To explore how this affects our estimates, we repeat the analysis shown in Table 5 using random subsamples of our data. Specifically, we draw ten 1/3 random subsamples of the birth cohort 1981-1989, to then repeat our analysis in each subsample. We then report the average coefficients and standard errors across those 10 subsamples. Table C.5 in the Appendix reports the results. The general pattern remains very similar: we still find that variation in $\hat{\lambda}_{r}$ rather than variation in $\hat{\rho}_{r}$ explains most of the variation in mobility across municipalities.
Second, we use our measures of average parental and grandparental years of schooling and earnings to perform the same analysis. As mentioned in Section 3, such measures could provide a better proxy for a family’s social status in each generation, and also reduce the influence of measurement and sampling error. The resulting estimates of $\lambda_{r}$ are similar to the estimates we find when considering individual parents and grandparents (as in Table 4). However, $\hat{\lambda}_{r}$ based on generation-specific averages render less spatial variation. That the average levels of $\hat{\lambda}_{r}$ are not very affected is not surprising, given that this adjustment approach affects $\beta_{-1,r}$ and $\beta_{-2,r}$ in a proportionally similar way, and $\hat{\lambda}_{r}=\hat{\beta}_{-2,r}/\hat{\beta}_{-1,r}$.
Location. In our baseline analysis, we allocate each family (lineage) to the child’s location at age 16. This is problematic as different generations from the same lineage may live in different municipalities. There is not a simple way to address this problem short of modeling the geographic mobility process as well, which is likely to related to intergenerational mobility process. However, we can explore whether our conclusions are sensitive to alternative definitions for the location of each family. In Appendix Table C.6 we repeat the analysis shown in Table 5, but use the mother’s location when she was young rather than the child’s location at age 16 to allocate each family to municipalities; see Section 2 for details. The overall pattern remains very similar as in our baseline. Again, variation in $\hat{\lambda}$ rather than $\hat{\rho}$ explains most of the variation in multigenerational correlations across regions. We therefore conclude that geographic mobility, while not explicitly accounted for in our analysis, is unlikely to affect our main conclusions.
7 Conclusions
Intergenerational mobility varies substantially across regions, and this variability may be informative about the mechanisms by which advantages are passed from parents to children. In this paper, we studied how multigenerational correlations vary across municipalities in Sweden, and argued that these multigenerational patterns provide additional insights on why parent-child correlations differ across regions. Our analysis therefore aimed to combine the advantages of two recent strands of the literature on geographic (e.g., CHKS_Where_QJE) and multigenerational patterns in mobility (e.g., LindahlPalme2014_IGE4Generations). We also linked to a recent debate on the relation between mobility and cross-sectional inequality (e.g., durlauf2022great).
As a starting point, we described the extent of regional variations in inter- and multigenerational educational and earnings mobility across Swedish municipalities, and their spatial patterns. Educational mobility tends to be lower in the more densely populated southern parts of Sweden and around major cities, while the pattern of earnings mobility is more dispersed. These regional rankings are largely stable to the choice of mobility measures. Different measures of educational mobility are highly correlated, and similar for inter- (parent-child) and multigenerational (grandparent-child) correlations along both the paternal and maternal lines. Regions also rank similarly across inter- and multigenerational correlations in earnings ranks, although these patterns are more dispersed. Moreover, areas with high earnings mobility are not necessarily those with high educational mobility.
These regional variations are interesting from a descriptive perspective, but our main purpose was to exploit them to learn more about intergenerational processes. We found that regional grandfather-child correlations are consistently larger than the square of their parent-child counterparts, implying that the latter only capture part of the intergenerational process, and tend to understate status transmission in the “long run” (across many generations). But while understating the level of transmission, parent-child correlations still capture differences in long-run transmission across regions. Conventional mobility statistics thus remain useful for comparative purposes, irrespectively of whether our interest centers on two- or multigenerational processes.
We further replicate the well-established pattern that regions characterized by high earnings inequality also tend to be characterized by low mobility, including across three generations. This relationship holds for most mobility measures that we considered, but the magnitudes vary. One interesting observation is that the inequality-mobility relation is particularly pronounced for educational mobility, especially for multigenerational correlations in education.
Finally, we illustrated that a comparison of inter- and multigenerational correlations can provide insights into how and why mobility differs across regions. Interpreting those regional differences through the lens of a latent factor model, we showed that they appear to be due to variation in the transmission of latent advantages from parents to children, rather than from differences in how those latent advantages convert into observable outcomes. A similar argument can be made with regard to the determinants of cross-sectional inequality. While these findings are stylized, they illustrate the potential value of integrating regional and multigenerational perspectives on intergenerational mobility.
Our analysis is subject to some important limitations. One such limitation is that we abstract from the role of geographic mobility, by allocating each family to the child’s municipality of residence. We provided evidence that our main conclusions are not sensitive to this choice, but given that geographical and intergenerational mobility processes are intertwined, it would be useful to account for migration patterns more explicitly in future work. Another important limitation is that parts of our analysis relied on steady-state assumptions, which are even less plausible in a multigenerational than an intergenerational context. Future research could aim to relax these constraints.
Data and Code Availability Statement
The paper relies on restricted-access individual-level data from Sweden. Due to the sensitive nature of these data, they are confidential and protected according to 24 ch. 8 ß of the Public Access to Information and Secrecy Act (2009:400). We are therefore prohibited from posting the data online. However, codes replicating the tables and figures in this article can be found on OpenICPSR, https://doi.org/10.3886/E226601V1. While we cannot share the microdata, the replication package includes all codes used to generate the results and instructions on how to apply for access to the data.
Online Appendix
Appendix A Supplementary Evidence for Sections 2 and 3
Figure A.1: Cumulative distribution of weights across municipalities
(a) Education correlations
(b) Earnings rank slopes
Notes: The figures show the distribution of weights (i.e. municipality-specific sample size) over municipalities, ordered from the smallest to the largest (from left to right) separately for education correlations (sub-figure a) and earnings rank slopes (sub-figure b). Each figure shows the CDF separately for the samples underlying the relevant intergenerational (father-child) and multigenerational (paternal grandfather-child) analyses.
Figure A.2: Inter- and multigenerational earnings rank slopes
Notes: Density of the estimated inter- and multigenerational coefficients in earnings ranks for parent-child and grandparent-child pairs across 290 Swedish municipalities (bandwidth: 0.01), based on birth cohorts 1981-1989. Observations are weighted by the number of parent-child pairs.
Figure A.3: Tests of whether $\triangle>0$
(a) Education
(b) Earnings ranks
Figure A.4: Inter- and multigenerational earnings mobility across Swedish municipalities
(a) Father-child earnings rank slope
(b) Grandfather-child earnings rank slope
Notes: The figures plot estimates of the father-child earnings rank slope (sub-figure a) and grandfather-child earnings rank slope (sub-figure b), across Swedish municipalities. The estimates are based on birth cohorts 1981-89.
Figure A.5: Intergenerational earnings mobility across Swedish municipalities, alternative measures
(a) Father-child earnings IGE
(b) Father-child absolute upward mobility
Notes: The figures plot estimates of the father-child intergenerational elasticity (IGE, sub-figure a) and absolute upward mobility, measured as the expected earnings rank of those with fathers at the 25th percentile of the earnings distribution (sub-figure b), across Swedish municipalities. The estimates are based on birth cohorts 1981-89.
Table A.1: Earnings measures
Child: Baseline/predicted Age 30-32 Age 30-39
Mean SD Mean SD Mean SD
Father:
     Baseline/predicted .211 .037 .158 .036 .169 .035
     5-year average .192 .040 .143 .036 .153 .036
Grandfather (pat.)
     Baseline/predicted .099 .064 .073 .061 .078 .059
     5-year average .065 .058 .053 .059 .056 .054
Note: Intergenerational and multigenerational rank coefficients for various earnings measures. Means and standard deviations across municipalities. For the child, columns 1-2 use our baseline measure with predicted earnings at age 40, columns 3-4 an average over ages 30-32, and columns 5-6 an average over ages 30-39. Note that only the earliest cohorts are observed up until age 39. We then use two separate measures for fathers and paternal grandfathers. Rows 1 and 3 use the baseline with predicted earnings at age 40 and rows 2 and 4 use a simple 5-year average around age 14 of the child (or of the father in the case of the grandfather). The underlying observation numbers differ across cells.
Table A.2: Inter- and multigenerational correlations by child gender
Sons Daughters
Yrs. education Earnings ranks Yrs. education Earnings ranks
Mean SD Mean SD Mean SD Mean SD
Father-child .327 .053 .238 .037 .270 .044 .207 .038
Mother-child .319 .047 .187 .036 .303 .036 .217 .033
Parental average-child .380 .050 .322 .043 .338 .039 .321 .045
Grandfather(pat.)-child .139 .054 .086 .038 .112 .045 .107 .036
Grandmother(pat.)-child .098 .045 .082 .040
Grandfather(mat.)-child .134 .049 .092 .038 .106 .046 .110 .037
Grandmother(mat.)-child .094 .043 .079 .036
Grandparental avg.-child .163 .055 .128 .038 .132 .046 .164 .063
Weighted by size Yes Yes Yes Yes Yes Yes Yes Yes
Note: Means and standard deviations across municipalities for different inter- and multigenerational measures, separately for sons and daughters. We do not report earnings rank coefficients for grandmothers due to smaller samples. However, the grandparental average earnings rank includes all four grandparents. All regressions are weighted by municipality-specific sample size in the main pooled samples for education and earnings ranks.
Figure A.6: Non-linear intergenerational relationship in education, national level
Notes: Binned scatter plot of the expectation of child years of education conditional on father’s years of education on the national level. Note that only 4% of observations are in the lowest bin.
Figure A.7: Non-linear multigenerational earnings rank slopes
(a) National level
(b) Separately for the five largest cities
Notes: Binned scatter plot of the expectation of child earnings rank conditional on grandfather’s earnings rank on the national level (sub-figure a) or municipal level (sub-figure b). Ranks are defined relative to the national distribution.
Figure A.8: Intergenerational earnings rank slopes by municipality size
Notes: Density plot of the estimated intergenerational coefficients in earnings ranks for father-child pairs across 163 “small” municipalities ($\leq$ 2,000 observations), 127 “large” municipalities (>> 2,000 observations), and similar numbers of small and large “placebo” municipalities (random allocation of observations across municipalities).
[…]
Appendix B Supplementary Evidence for Sections 4 and 5 Table B.1: Correlation matrix
$\hat{\beta}_{f}^{edu}$ $\hat{\beta}_{fgf}^{edu}$ $\hat{\beta}_{m}^{edu}$ $\hat{\beta}_{mgf}^{edu}$ $\hat{\beta}_{f}^{rank}$ $\hat{\beta}_{fgf}^{rank}$ $\hat{\beta}_{m}^{rank}$ $\hat{\beta}_{mgf}^{rank}$ $\hat{\beta}_{f}^{IGE}$ $\hat{\beta}_{fgf}^{IGE}$ $\hat{\beta}_{m}^{IGE}$ $\hat{\beta}_{mgf}^{IGE}$ $\hat{\beta}_{f}^{edu}$ 1.00 $\hat{\beta}_{fgf}^{edu}$ 0.79 1.00 $\hat{\beta}_{m}^{edu}$ 0.83 0.69 1.00 $\hat{\beta}_{mgf}^{edu}$ 0.75 0.80 0.72 1.00 $\hat{\beta}_{f}^{rank}$ 0.06 -0.00 0.08 -0.01 1.00 $\hat{\beta}_{fgf}^{rank}$ 0.35 0.33 0.28 0.31 0.51 1.00 $\hat{\beta}_{m}^{rank}$ 0.05 0.04 0.18 0.09 0.57 0.35 1.00 $\hat{\beta}_{mgf}^{rank}$ 0.29 0.25 0.30 0.28 0.51 0.56 0.44 1.00 $\hat{\beta}_{f}^{IGE}$ -0.11 -0.21 -0.08 -0.20 0.91 0.39 0.49 0.38 1.00 $\hat{\beta}_{fgf}^{IGE}$ 0.23 0.19 0.17 0.17 0.54 0.94 0.39 0.54 0.47 1.00 $\hat{\beta}_{m}^{IGE}$ -0.03 -0.07 0.10 -0.02 0.55 0.29 0.95 0.36 0.55 0.35 1.00 $\hat{\beta}_{mgf}^{IGE}$ 0.14 0.08 0.16 0.12 0.55 0.53 0.48 0.94 0.47 0.56 0.43 1.00
Notes: Correlations between different estimates of municipality-specific inter- and multigenerational measures. We exclude municipalities with fewer than 1000 observations and weight the correlation by sample size. Subscript $f$ stands for fathers (e.g., $\hat{\beta}_{f}^{edu}$ is the father-child correlation in years of schooling), $m$ for mothers, fgffgf for paternal grandfathers, mgfmgf for maternal grandfathers.
Figure B.1: Multigenerational Great Gatsby Curve, Alternative Inequality Measure
Notes: The figure plots the multigenerational (child vs. paternal grandfather) rank slope in earnings against earnings inequality among the fathers. Scatter plots and linear fit restricted to municipalities for which at least 1,000 father-child pairs are observed. The linear regression line is weighted by the number of father-child pairs.
[…]
Appendix C Supplementary Evidence for Section 6 Table C.1: Regional variation in the latent factor model, sons and daughters
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$ Panel A: Sons Years of schooling Earnings ranks (1) (2) (3) (4) (5) (6) $\hat{\rho}_{r}$ (“returns”) -0.176*** 0.086*** -0.010* -0.001 (0.023) (0.023) (0.004) (0.000) $\hat{\lambda}_{r}$ (“transferability”) 0.335*** 0.476*** 0.193*** 0.195*** (0.016) (0.027) (0.009) (0.012) # Municipalities 275 290 275 272 289 272 Adj. $R^{2}$ 0.386 0.849 0.901 0.084 0.811 0.762 Panel B: Daughters Dependent variable: Grandfather-child correlation $\beta_{-2,r}$ Years of schooling Earnings ranks (1) (2) (3) (4) (5) (6) $\hat{\rho}_{r}$ (“returns”) -0.130*** 0.049** -0.080*** 0.008 (0.013) (0.017) (0.016) (0.010) $\hat{\lambda}_{r}$ (“transferability”) 0.270*** 0.369*** 0.167*** 0.173*** (0.012) (0.033) (0.009) (0.015) # Municipalities 268 290 268 288 289 287 Adj. $R^{2}$ 0.399 0.823 0.821 0.272 0.665 0.656
Notes: The dependent variable is the grandfather-child correlation in years of schooling (columns 1-3) or in earnings ranks (columns 4-6) across 290 Swedish municipalities, separately for sons (Panel A) and daughters (Panel B). The independent variables are estimates based on the corresponding parent-child and grandparent-child correlations, see equations (6) and (7). Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.001.
Table C.3: Regional variation in the latent factor model, balanced sample
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$
Years of schooling Earnings ranks
(1) (2) (3) (4) (5) (6)
$\hat{\rho}$ (“returns”) -0.086*** 0.034*** -0.046* 0.013*
(0.023) (0.014) (0.019) (0.012)
$\hat{\lambda}$ (“transferability”) 0.363*** 0.446*** 0.193*** 0.205***
(0.016) (0.022) (0.008) (0.015)
# Municipalities 280 290 280 283 289 282
AR2 0.234 0.847 0.878 0.137 0.695 0.665
Notes: The dependent variable is the grandfather-child correlation in years of schooling (columns 1-3) or the grandfather-child correlation in earnings ranks (columns 4-6) across 290 Swedish municipalities. The independent variables are estimates based on the corresponding father-child and grandfather-child correlations, see equations (6) and (7). All correlations are estimated on a balanced sample with complete child-father-grandfather lineages. Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.005.
Table C.4: Regional variation in the latent factor model, binary education variable
Dependent variable: $\beta_{-2,r}$
High/low schooling
(1) (2) (3)
$\hat{\rho}$ (“returns”) -0.022* 0.002
(0.011) (0.002)
$\hat{\lambda}$ (“transferability”) 0.212*** 0.229***
(0.007) (0.012)
# Municipalities 271 290 271
AR2 0.191 0.880 0.856
Notes: The dependent variable is the grandfather-child correlation in an indicator for high/low schooling, defined relative to the median of the respective generation across 290 Swedish municipalities. The independent variables are estimates based on the corresponding father-child and grandfather-child correlations, see equations (6) and (7). Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Standard errors are clustered at the level of 25 regions:* p<0.05, ** p<0.01, *** p<0.005.
Table C.5: Regional variation in the latent factor model, log-linear model
Dependent variable: Log grandfather-child correlation $log(\beta_{-2,r})$
Years of schooling Earnings ranks
(1) (2) (3) (4) (5) (6)
[…]
(0.119) (0.000) (0.163) (0.000)
[…]
(0.028) (0.000) (0.026) (0.000)
# Municipalities 280 280 280 283 283 283
AR2 0.645 0.938 1.000 0.556 0.886 1.000
[…]
Table C.6: Regional variation and the latent factor model, average across sub-samples
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$
Years of schooling Earnings ranks
(1) (2) (3) (4) (5) (6)
$\hat{\rho}_{r}$ (“returns”) -0.075*** 0.013 -0.079*** 0.001**
(0.019) (0.011) (0.014) (0.012)
$\hat{\lambda}_{r}$ (“transferability”) 0.272*** 0.310 0.168*** 0.164***
(0.022) (0.041) (0.014) (0.025)
# Municipalities 271 289 270 270 288 269
Adj. $R^{2}$ 0.247 0.767 0.741 0.294 0.716 0.647
Notes: The table reports the average coefficient estimates and standard errors across 10 random subsamples of size 1/3. The dependent variable is the grandfather-child correlation in years of schooling (columns 1-3) or the grandfather-child correlation in earnings ranks (columns 4-6) across 290 Swedish municipalities. The independent variables are estimates based on the corresponding father-child and grandfather-child correlations, see equations (6) and (6). Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.001.
Table C.7: Regional variation in the latent factor model, mother’s location
Dependent variable: Grandfather-child correlation $\beta_{-2,r}$
Years of schooling Earnings ranks
(1) (2) (3) (4) (5) (6)
$\hat{\rho}$ (“returns”) -0.220*** 0.124*** -0.117*** 0.051
(0.026) (0.030) (0.035) (0.032)
$\hat{\lambda}$ (“transferability”) 0.382*** 0.511*** 0.250*** 0.301***
(0.013) (0.029) (0.008) (0.029)
# Municipalities 280 281 280 278 281 278
AR2 0.445 0.889 0.932 0.300 0.831 0.852
Notes: The dependent variable is the grandfather-child correlation in years of schooling (columns 1-3) or the grandfather-child correlation in earnings ranks (columns 4-6) across 290 Swedish municipalities. The independent variables are estimates based on the corresponding father-child and grandfather-child correlations, see equations (6) and (7). All correlations are estimated based on the mother’s location. Birth cohorts 1981-89, weighted by the number of grandfather-child pairs in each municipality. Robust standard errors in parentheses:* p<0.05, ** p<0.01, *** p<0.005.